# Federated Learning: A new frontier in the exploration of multi-institutional medical imaging data

### v1

@@ 6.2 Data heterogeneity
Specifically, the Inside and Outside model Personalization in FL (IOP-FL) framework [Jiang et al., 2023] employs a locally adapted model for inside personalization (i.e., clients involved in the joint training of the global model) and applies test-time routing for outside personalization (i.e., for noninvolved clients), enabling institutions that are not part of the FL process to use the trained model effectively. By combining local gradients (client-specific optimization) and global gradients (common knowledge) during local model training, the framework generates a model that can learn the common pattern and fit the local data distribution. The IOP-FL approach allows for a flexible combination of training knowledge and information from test data. The technique has been used for prostate segmentation from MRI data and optic disc and cup segmentation from retinal fundus images, demonstrating its advantages over traditional FL algorithms like FedAvg or FedBN.
@@
The Split Averaging (SplitAVG) by Zhang et al. [2022a] (see Fig. 10A) divides a DL network into two sub-networks, respectively, local and global at a specific cut layer. Model training involves forward and backpropagation. First, intermediate feature maps are obtained from the local sub-network and sent with corresponding labels for concatenation to the central server. These concatenated feature maps are then propagated into the global sub-network. Second, gradients are backpropagated from the last layer to the first layer of the global model, with the first global layer gradients transferred back to all clients for further backpropagation. Finally, each client and the server update their sub-networks. An improved version of the SplitAVG technique is SplitAVG-v2 [Zhang et al., 2022a], which allows the preservation of locally confidential labels without sharing them with the server. Both methods have been used effectively for various medical imaging tasks, including the binary classification of diabetic retinopathy, the prediction of bone age, and the segmentation of brain tumors.
Specificity-preserving FL algorithm for MR Image Reconstruction (FedMRI) [Feng et al., 2023] is another method based on dividing the model into two parts: a globally shared encoder and a client-specific decoder (see Fig. 10B). In addition to the standard loss function, the FedMRI employs an extra specialized weighted contrastive regularization loss that enables further correction for local updates and provides the model with global recognition ability.
$\mathcal{L}_{\mathrm{con}_{c}}(\mathbf{\Theta}_{E}^{r},\mathbf{\Theta}_{E_{c}}^{r},\{\mathbf{\Theta}_{E_{c^{\prime}}}^{r-1}|c^{\prime}\in C\})=\frac{||\mathbf{\Theta}_{E_{c}}^{r}-\mathbf{\Theta}_{E}^{r}||_{1}}{\sum_{c^{\prime}=1}^{C}||\mathbf{\Theta}_{E_{c^{\prime}}}^{r-1}-\mathbf{\Theta}_{E_{c}}^{r}||_{1}}$ (33)
with $\mathbf{\Theta}_{E}^{r}$ and $\mathbf{\Theta}_{E_{c}}^{r}$ being the parameters of global and local encoders at round $r$. FedMRI has already shown an improved accuracy in the reconstruction of MRI data compared to traditional FL methods such as FedAvg, FedBN, and FedProx
The Federated Learning of Generative Image Priors (FedGIMP) [Elmas et al., 2023] employs an unconditional adversarial model divided into a shared generator and a local discriminator (see Fig. 10C). After every epoch, the local generators are sent to the server and aggregated similarly to the FedAvg method. A model trained using this technique can be employed for inference, allowing it to adapt to a new dataset not included during training. During the inference, the trained MRI prior, which is represented as an adversarial model capable of synthesizing MRI images using site-specific latent variables, is combined with a subject-specific imaging operator and adjusted to minimize data consistency loss. FedGIMP has shown a notable improvement compared to other FL methods in the reconstruction of MRI data.
Another method based on split learning is the Personalized FL Method for MRI Syntheses (pFLSynth) by Dalmaz et al. [2024]. Here, the adversarial model is divided into a globally shared downstream synthesizer and mapper and a client-specific upstream synthesizer and discriminator (see Fig. 10D). Finally, only the downstream synthesizer and mapper parameters are aggregated on the central server
@@
where $\mathbf{\Theta}_{\mathrm{syn}}^{r+1}$/$\mathbf{\Theta}_{\mathrm{map}}^{r+1}$ are the aggregated parameters of the downstream synthesizer/mapper in round $(r+1)$ and $\mathbf{\Theta}^{r}_{\mathrm{syn}_{c}}$/$\mathbf{\Theta}^{r}_{\mathrm{map}_{c}}$ are client-specific upstream synthesizer/mapper parameters. The pFLSynth method has demonstrated its capabilities in synthesizing multi-contrast MRI images.
@@
$\mathcal{L}_{\mathrm{gwc}_{c}}(\mathbf{\Theta}_{\mathrm{den}}^{r},\mathbf{\Theta}_{\mathrm{den}_{c}}^{r})=||\mathbf{\Theta}_{\mathrm{den}}^{r}-\mathbf{\Theta}_{\mathrm{den}_{c}}^{r}||_{2}^{2}$ (35)
with $\mathbf{\Theta}_{\mathrm{den}}^{r}$ and $\mathbf{\Theta}_{\mathrm{den}_{c}}^{r}$ being the parameters of the global, averaged by the central server, and local denoising network, respectively. FedFTN has been used successfully to denoise low-count PET images.
In addition to dividing the model parameters into local and global parts, it is also possible to disentangle the parameter space into shape and appearance and share only the shape parameter between the clients. This separation uses shared structural information from various clients and reduces domain shifts by enabling personalized appearance models. This kind of approach is used by Federated Disentanglement (FedDis) algorithm [Bercea et al., 2022]. The FedDis assumes that the anatomical structure of the data representing the brain is similar across all collaborators and that the heterogeneity is mainly due to differences in data collection devices and their parameters. The approach has been effectively applied in segmenting the MRI data, exceeding the accuracy of locally trained models.
Data sharing: The next group of methods handling data heterogeneity follows data sharing scheme, which can be direct [Guo et al., 2021] or indirect [Liu et al., 2021a, Luo and Wu, 2022, Zhao and Huang, 2023]. Direct data sharing involves distributing the publicly available dataset among all clients, while indirect data sharing indicates that clients exclusively exchange details about the data, such as its distribution, instead of the actual data (see Fig. 11 for a graphical explanation).
The first indirect data sharing method mentioned here is the Episodic Learning in Continuous Frequency Space (ELCFS) by Liu et al. [2021a], which enables multiple clients to benefit from data distributions from different sources. The ELCFS employs a continuous frequency space interpolation mechanism to share distribution information between clients. Initially, low-level distributions, such as style, are extracted from the frequency domain of local participant data as an amplitude spectrum. These distributions are then compiled into a centralized distribution bank that is accessible to all clients. Consequently, this approach enables the generation of images with modified appearances that reflect the distribution characteristics of other clients. Additionally, ELCFS employs an episodic meta-learning approach that learns generalizable model parameters by simulating train and test domain shifts. Combined with boundary-oriented meta-optimization, which focuses on distinguishing boundary-related features from background-related ones, it enhances domain generalization. This method has been successfully used to segment the optic disc and cup in images of the retinal fundus and for prostate segmentation in $T_{2}$-weighted MRI scans.
@@
Another example of an indirect data sharing method is Federated Learning with Shared Label Distribution (FedSLD) [Luo and Wu, 2022]. This method assumes that the number of samples from every class for all clients is known. This information is then used in the local objective function to ensure that each class contributes proportionally, relative to the total number of samples from that class across all clients within the FL architecture. FedSLD has been evaluated for the classification of liver and colorectal cancer images, achieving better results than FedAvg or FedProx.
Yet another indirect method called Distribution Information Sharing Federated Learning approach (FedDIS) by Zhao and Huang [2023] enables the involved clients to share their unique local data distributions with each other to create an independent and identically distributed (IID) dataset. The method trains a variational autoencoder (VAE) and subsequently maps local medical data to a hidden space via the encoder. The distributional information about the mapped data in the hidden space is then estimated and shared among the clients. Clients can generate a new dataset based on the received distribution information using the VAE decoder. FedDIS has significantly improved the accuracy of diagnosing Alzheimer’s disease from MRI data in a non-IID scenario.
@@
The first of two domain adaptation approaches introduced here is the Multi-site fMRI Analysis Framework [Li et al., 2020b]. The technique uses two domain adaptation methods: a mixture of experts (MoE) and adversarial domain alignment. The MoE is a gating network that dynamically assigns weight values during the global aggregation of local models. This approach is used for adaptation near the output layer. On the other hand, adversarial domain alignment is a technique employed for adaptation on the data knowledge representation level. The method of Li et al. [2020b] has been successfully used to classify data from individuals with autism.
The second method, previously mentioned under the umbrella of data sharing techniques, is the FL-MR with Cross-site Modeling (FL-MRCM) [Guo et al., 2021]. The FL-MRCM aims to solve the domain shift problem by using the data of the target site, along with the adversarial domain identifier, and the reconstruction network consisting of an encoder and a decoder. The encoder projects the input from the source and target sites into the latent space, and the identifier aligns the distribution of the latent space between each pair of source-target domains. FL-MRCM has been used successfully for MRI reconstruction, achieving results comparable to those obtained from centrally trained models.
@@
The Federated Conditional Mutual Learning (FedCM) by Huang et al. [2021] uses the local performance of the clients and the similarity to improve the overall accuracy of the global model. The client performance is evaluated based on the cross-entropy loss calculated on their private test set, while similarity is assessed using predicted logits (class scores) from public data. Each client periodically uploads the loss value and the predicted logits to the server. The server then sends back the set of losses and logits of the remaining clients (see Fig. 12). The client updates its knowledge by mutual distillation and fine-tuning its local dataset for personalization. FedCM has demonstrated superior performance in classifying Alzheimer’s disease using $T_{1}$-weighted MRI data, outperforming transfer learning and FedAvg approaches.
@@
Next, FedZaCt employs a cross-teaching method, which enables local clients to use the Z-average models to train their models. To achieve this, an additional component is included in the overall loss function. This component measures the difference between the predictions given by the local model and those of the Z-average models. FedZaCt has been successfully used to segment various medical images, including dermatoscopic images with pigmented skin lesions and CT images showing the aortic vessel region.
The Privacy-preserving Generative Adversarial Network (MixFedGAN) by Yang et al. [2023] incorporates knowledge distillation into the local model. The method introduces a distillation regularization loss, which employs a Kullback-Leibler divergence to minimize the difference between the local and global models. MixFedGAN has been evaluated effectively on heterogeneous datasets of COVID-19 infection and prostate MRI, demonstrating its advantages over other FL methods, including FedAvg, FedProx, and FedBN.
@@
The FedIIC [Wu et al., 2023a] is an example of a contrastive learning-based approach to address the challenges of class imbalance. FedIIC combines feature and classifier learning. Firstly, feature learning covers two levels of contrastive learning: intra- and inter-client. On the one hand, intra-client contrastive learning extracts class-specific features by contrasting representations of images within the same class against those from different classes. On the other hand, inter-client contrastive learning involves contrasting features across different clients. This procedure employs class-wise prototypes, which are generated from the global model weights and shared among all clients. By aligning features from various clients, the model achieves better generalization. Secondly, classifier learning enables the dynamic adjustment of the classifier in response to real-time difficulty. These difficulties are assessed by the average cross-entropy loss of all samples within a specific class using a global model. Such an adjustment helps to prevent the model from becoming biased towards majority classes, ensuring fair performance across all classes. FedIIC has demonstrated significant advantages for the classification of dermatoscopic and CT images, compared to standard techniques such as FedAvg or FedProx.
Synthesizing additional data: This approach has been designed to deal with non-IID data and involves synthesizing additional virtual data. The Federated Medical Image Analysis with Virtual Sample Synthesis [Zhu and Luo, 2022] synthesizes training samples using local and global models. To improve the generalization of the local model, virtual adversarial training is applied. Unlike adversarial training, this method generates adversarial neighbors that closely resemble the selected input sample in terms of raw pixels but produce different predictions. In addition, the global model is used to generate high-confident samples, which enables the alignment of the local distribution to the global one. This method has been successfully used to detect and classify histological data from various parts of the body.
Replacing the architecture: The Federated Learning with Hyper‑network [Lin et al., 2023] enables hyper-network training instead of global model training. The hyper-network is a neural network that produces parameters for a larger target model. The hyper-network uses a set of inputs that contain information about the structure of local model parameters and generates parameters for each layer of the target network. This approach has been successfully applied in classification.
@@ 6.3 Label deficiency
In Federated Contrastive Learning (FCL) framework [Wu et al., 2021], the clients exchange features with each other during the pre-training procedure to provide more diverse data. Local contrastive learning involves using two encoders: the main encoder, which is learned and used for initialization during fine-tuning, and the momentum encoder, a slowly evolving version of the main encoder, generating features to discriminate and intended for sharing. Then, the FCL aligns similar features among collaborators using the structural similarity of the data. FCL has been designed for the volumetric segmentation of medical images and evaluated on a cardiac MRI dataset, demonstrating its advantages over other techniques that employ contrastive learning.
The Robust Federated Contrastive Learning (FedMoCo) framework [Dong and Voiculescu, 2021] combines contrastive learning and metadata transfer, facilitating the augmentation of inter-node statistical data. During contrastive learning, the data augmentation procedure creates positive pairs by generating two random versions of the same image through various transformations, such as flipping and rotating. In contrast, negative pairs are formed using different versions of two distinct images. Eventually, the metadata transfer involves transferring the statistical information of the transformed features between clients. FedMoCo has been used effectively to detect COVID-19 from chest X-ray data.
@@
FATE Comprehensive framework Python Secure computation protocols, secure multi-party computation, task security scheduling, homomorphic encryption https://fate.fedai.org/
@@
Federated Alternate Training (FAT) [Mushtaq et al., 2023] alters the training procedure between annotated and unannotated data. On the one hand, the method uses labeled datasets to fine-tune the global model parameters. On the other hand, clients with unlabeled datasets use the global model as a target model to generate pseudo-labels for self-supervised learning. Model aggregation is performed alternately. First, the global model parameters are aggregated for several rounds for clients with annotated data. Then, for the following few rounds, the aggregation is carried out for clients with unannotated data. FAT has been used in medical image segmentation and has been tested on CT and MRI data.
In Class Imbalanced Medical Image Classification Based on Semi-Supervised Federated Learning [Liu et al., 2023b], local clients do not have labels at all, while the labeled data is sent from the central server to each collaborator, allowing them to participate in semi-supervised training. Each client combines the labeled and unlabeled data to construct an extended dataset used to generate pseudo-labels. This method considers the class imbalance factor, which is determined based on annotated data. The technique has been successfully used to classify dermatoscopic images, outperforming FedAvg and other semi-supervised learning techniques.
Generating synthetic data: Federated Graph Learning with Network Inpainting for Population-Based Disease Prediction (FedNI) [Peng et al., 2023] synthesizes data for graph convolutional neural networks (GCN). This method employs a network inpainting module that includes a missing node generator. It aims to predict missing nodes and augment edge connections. As a result, a larger fused local graph is generated, which serves as input to the GCN node classifier. FendNI has been used to predict neurological disorders and diseases using MRI data.
Knowledge distillation: Federated Learning with Knowledge Distillation for Multi-organ Segmentation with Partially Labeled Datasets [Kim et al., 2024] involves global and local versions of knowledge distillation. The global knowledge distillation is designed to prevent catastrophic forgetting by preserving predictions for organs with missing labels. In this context, catastrophic forgetting refers to the loss of knowledge about the segmentation of specific organs during local training, which occurs once clients own only partially annotated datasets. In local knowledge distillation, each client trains a model for a specific organ before starting the FL procedure and shares its parameters with collaborators. The client uses predictions from another participant’s randomly selected organ-specific model during the local training. This approach is especially beneficial for multi-organ segmentation, where each client has labels for only one organ [Kim et al., 2024].

### v2

@@ 6.2 Data heterogeneity
Specifically, the Inside and Outside model Personalization in FL (IOP-FL) framework [Jiang et al., 2023] employs a locally adapted model for inside personalization (i.e., clients involved in the joint training of the global model) and applies test-time routing for outside personalization (i.e., for noninvolved clients), enabling institutions that are not part of the FL process to use the trained model effectively. By combining local gradients (client-specific optimization) and global gradients (common knowledge) during local model training, the framework generates a model that can learn the common pattern and fit the local data distribution. The IOP-FL approach allows for a flexible combination of training knowledge and information from test data.
@@
The Split Averaging (SplitAVG) by Zhang et al. [2022a] (see Fig. 11A) divides a DL network into two sub-networks, respectively, local and global at a specific cut layer. Model training involves forward and backpropagation. First, intermediate feature maps are obtained from the local sub-network and sent with corresponding labels for concatenation to the central server. These concatenated feature maps are then propagated into the global sub-network. Second, gradients are backpropagated from the last layer to the first layer of the global model, with the first global layer gradients transferred back to all clients for further backpropagation. Finally, each client and the server update their sub-networks. An improved version of the SplitAVG technique is SplitAVG-v2 [Zhang et al., 2022a], which allows the preservation of locally confidential labels without sharing them with the server. Specificity-preserving FL algorithm for MR Image Reconstruction (FedMRI) [Feng et al., 2023] is another method based on dividing the model into two parts: a globally shared encoder and a client-specific decoder (see Fig. 11B). In addition to the standard loss function, the FedMRI employs an extra specialized weighted contrastive regularization loss that enables further correction for local updates and provides the model with global recognition ability.
[…]
with $\mathbf{\Theta}_{E}^{r}$ and $\mathbf{\Theta}_{E_{c}}^{r}$ being the parameters of global and local encoders at round $r$.
The Federated Learning of Generative Image Priors (FedGIMP) [Elmas et al., 2023] employs an unconditional adversarial model divided into a shared generator and a local discriminator (see Fig. 11C). After every epoch, the local generators are sent to the server and aggregated similarly to the FedAvg method. A model trained using this technique can be employed for inference, allowing it to adapt to a new dataset not included during training. During the inference, the trained MRI prior, which is represented as an adversarial model capable of synthesizing MRI images using site-specific latent variables, is combined with a subject-specific imaging operator and adjusted to minimize data consistency loss. Another method based on split learning is the Personalized FL Method for MRI Syntheses (pFLSynth) by Dalmaz et al. [2024]. Here, the adversarial model is divided into a globally shared downstream synthesizer and mapper and a client-specific upstream synthesizer and discriminator (see Fig. 11D). Finally, only the downstream synthesizer and mapper parameters are aggregated on the central server
@@
where $\mathbf{\Theta}_{\mathrm{syn}}^{r+1}$$\mathbf{\Theta}_{\mathrm{map}}^{r+1}$ are the aggregated parameters of the downstream synthesizer/mapper in round $(r+1)$ and $\mathbf{\Theta}^{r}_{\mathrm{syn}_{c}}$$\mathbf{\Theta}^{r}_{\mathrm{map}_{c}}$ are client-specific upstream synthesizer/mapper parameters.
@@
[…]
with $\mathbf{\Theta}_{\mathrm{den}}^{r}$ and $\mathbf{\Theta}_{\mathrm{den}_{c}}^{r}$ being the parameters of the global, averaged by the central server, and local denoising network, respectively.
In addition to dividing the model parameters into local and global parts, it is also possible to disentangle the parameter space into shape and appearance and share only the shape parameter between the clients. This separation uses shared structural information from various clients and reduces domain shifts by enabling personalized appearance models. This kind of approach is used by Federated Disentanglement (FedDis) algorithm [Bercea et al., 2022]. The FedDis assumes that the anatomical structure of the data representing the brain is similar across all collaborators and that the heterogeneity is mainly due to differences in data collection devices and their parameters.
Data sharing: The next group of methods handling data heterogeneity follows data sharing scheme, which can be direct [Guo et al., 2021] or indirect [Liu et al., 2021a, Luo and Wu, 2022, Zhao and Huang, 2023, Li et al., 2025a]. Direct data sharing involves distributing the publicly available dataset among all clients, while indirect data sharing indicates that clients exclusively exchange details about the data, such as its distribution, instead of the actual data (see Fig. 12 for a graphical explanation).
The first indirect data sharing method mentioned here is the Episodic Learning in Continuous Frequency Space (ELCFS) by Liu et al. [2021a], which enables multiple clients to benefit from data distributions from different sources. The ELCFS employs a continuous frequency space interpolation mechanism to share distribution information between clients. Initially, low-level distributions, such as style, are extracted from the frequency domain of local participant data as an amplitude spectrum. These distributions are then compiled into a centralized distribution bank that is accessible to all clients. Consequently, this approach enables the generation of images with modified appearances that reflect the distribution characteristics of other clients. Additionally, ELCFS employs an episodic meta-learning approach that learns generalizable model parameters by simulating train and test domain shifts. Combined with boundary-oriented meta-optimization, which focuses on distinguishing boundary-related features from background-related ones, it enhances domain generalization.
@@
Another example of an indirect data sharing method is Federated Learning with Shared Label Distribution (FedSLD) [Luo and Wu, 2022]. This method assumes that the number of samples from every class for all clients is known. This information is then used in the local objective function to ensure that each class contributes proportionally, relative to the total number of samples from that class across all clients within the FL architecture.
Yet another indirect method called Distribution Information Sharing Federated Learning approach (FedDIS) by Zhao and Huang [2023] enables the involved clients to share their unique local data distributions with each other to create an independent and identically distributed (IID) dataset. The method trains a variational autoencoder (VAE) and subsequently maps local medical data to a hidden space via the encoder. The distributional information about the mapped data in the hidden space is then estimated and shared among the clients. Clients can generate a new dataset based on the received distribution information using the VAE decoder.
Another method of indirect data sharing is Hepa-FedBoost [Li et al., 2025a]. Hepa-FedBoost sends anonymized data summaries, known as prototypes, along with model parameters to the server. Each client generates two types of prototypes during local training: 1) class-specific prototypes, computed as the mean feature vector across samples of a specific class, and 2) domain-agnostic prototypes, which aggregate global appearance cues across all data patches regardless of the class. These prototypes enable the central server to group clients with similar data characteristics. Client clusters are updated regularly, and model parameter aggregation occurs separately for each cluster. Hepa-FedBoost utilizes a technique called global-local contrastive boosting to distinguish between positive and negative samples. This approach helps refine decision boundaries, which are the lines, curves, or hyperplanes in a feature space that separate different classes.
@@
The first of two domain adaptation approaches introduced here is the Multi-site fMRI Analysis Framework [Li et al., 2020b]. The technique uses two domain adaptation methods: a mixture of experts (MoE) and adversarial domain alignment. The MoE is a gating network that dynamically assigns weight values during the global aggregation of local models. This approach is used for adaptation near the output layer. On the other hand, adversarial domain alignment is a technique employed for adaptation on the data knowledge representation level.
The second method, previously mentioned under the umbrella of data sharing techniques, is the FL-MR with Cross-site Modeling (FL-MRCM) [Guo et al., 2021]. The FL-MRCM aims to solve the domain shift problem by using the data of the target site, along with the adversarial domain identifier, and the reconstruction network consisting of an encoder and a decoder. The encoder projects the input from the source and target sites into the latent space, and the identifier aligns the distribution of the latent space between each pair of source-target domains.
@@
The Federated Conditional Mutual Learning (FedCM) by Huang et al. [2021] uses the local performance of the clients and the similarity to improve the overall accuracy of the global model. The client performance is evaluated based on the cross-entropy loss calculated on their private test set, while similarity is assessed using predicted logits (class scores) from public data. Each client periodically uploads the loss value and the predicted logits to the server. The server then sends back the set of losses and logits of the remaining clients (see Fig. 13). The client updates its knowledge by mutual distillation and fine-tuning its local dataset for personalization.
@@
Next, FedZaCt employs a cross-teaching method, which enables local clients to use the Z-average models to train their models. To achieve this, an additional component is included in the overall loss function. This component measures the difference between the predictions given by the local model and those of the Z-average models.
The Privacy-preserving Generative Adversarial Network (MixFedGAN) by Yang et al. [2023] incorporates knowledge distillation into the local model. The method introduces a distillation regularization loss, which employs a Kullback-Leibler divergence to minimize the difference between the local and global models.
@@
The FedIIC [Wu et al., 2023] is an example of a contrastive learning-based approach to address the challenges of class imbalance. FedIIC combines feature and classifier learning. Firstly, feature learning covers two levels of contrastive learning: intra- and inter-client. On the one hand, intra-client contrastive learning extracts class-specific features by contrasting representations of images within the same class against those from different classes. On the other hand, inter-client contrastive learning involves contrasting features across different clients. This procedure employs class-wise prototypes, which are generated from the global model weights and shared among all clients. By aligning features from various clients, the model achieves better generalization. Secondly, classifier learning enables the dynamic adjustment of the classifier in response to real-time difficulty. These difficulties are assessed by the average cross-entropy loss of all samples within a specific class using a global model. Such an adjustment helps to prevent the model from becoming biased towards majority classes, ensuring fair performance across all classes.
Synthesizing additional data: This approach has been designed to deal with non-IID data and involves synthesizing additional virtual data. The Federated Medical Image Analysis with Virtual Sample Synthesis [Zhu and Luo, 2022] synthesizes training samples using local and global models. To improve the generalization of the local model, virtual adversarial training is applied. Unlike adversarial training, this method generates adversarial neighbors that closely resemble the selected input sample in terms of raw pixels but produce different predictions. In addition, the global model is used to generate high-confident samples, which enables the alignment of the local distribution to the global one.
Replacing the architecture: The Federated Learning with Hyper‑network [Lin et al., 2023] enables hyper-network training instead of global model training. The hyper-network is a neural network that produces parameters for a larger target model. The hyper-network uses a set of inputs that contain information about the structure of local model parameters and generates parameters for each layer of the target network.
@@ 6.3 Label deficiency
In Federated Contrastive Learning (FCL) framework [Wu et al., 2021], the clients exchange features with each other during the pre-training procedure to provide more diverse data. Local contrastive learning involves using two encoders: the main encoder, which is learned and used for initialization during fine-tuning, and the momentum encoder, a slowly evolving version of the main encoder, generating features to discriminate and intended for sharing. Then, the FCL aligns similar features among collaborators using the structural similarity of the data.
The Robust Federated Contrastive Learning (FedMoCo) framework [Dong and Voiculescu, 2021] combines contrastive learning and metadata transfer, facilitating the augmentation of inter-node statistical data. During contrastive learning, the data augmentation procedure creates positive pairs by generating two random versions of the same image through various transformations, such as flipping and rotating. In contrast, negative pairs are formed using different versions of two distinct images. Eventually, the metadata transfer involves transferring the statistical information of the transformed features between clients.
@@
FATE Comprehensive framework Python Secure computation protocols, secure multi-party computation, task security scheduling, homomorphic encryption https://github.com/FederatedAI/FATE
@@
Fed-BioMed Comprehensive framework Python Secure aggregation, differential privacy https://fedbiomed.org/
@@
Federated Alternate Training (FAT) [Mushtaq et al., 2023] alters the training procedure between annotated and unannotated data. On the one hand, the method uses labeled datasets to fine-tune the global model parameters. On the other hand, clients with unlabeled datasets use the global model as a target model to generate pseudo-labels for self-supervised learning. Model aggregation is performed alternately. First, the global model parameters are aggregated for several rounds for clients with annotated data. Then, for the following few rounds, the aggregation is carried out for clients with unannotated data.
In Class Imbalanced Medical Image Classification Based on Semi-Supervised Federated Learning [Liu et al., 2023b], local clients do not have labels at all, while the labeled data is sent from the central server to each collaborator, allowing them to participate in semi-supervised training. Each client combines the labeled and unlabeled data to construct an extended dataset used to generate pseudo-labels. This method considers the class imbalance factor, which is determined based on annotated data.
Generative methods: Federated Graph Learning with Network Inpainting for Population-Based Disease Prediction (FedNI) [Peng et al., 2023] synthesizes data for graph convolutional neural networks (GCN). This method employs a network inpainting module that includes a missing node generator. It aims to predict missing nodes and augment edge connections. As a result, a larger fused local graph is generated, which serves as input to the GCN node classifier.
Knowledge distillation: Federated Learning with Knowledge Distillation for Multi-organ Segmentation with Partially Labeled Datasets [Kim et al., 2024] involves global and local versions of knowledge distillation. The global knowledge distillation is designed to prevent catastrophic forgetting by preserving predictions for organs with missing labels. In this context, catastrophic forgetting refers to the loss of knowledge about the segmentation of specific organs during local training, which occurs once clients own only partially annotated datasets. In local knowledge distillation, each client trains a model for a specific organ before starting the FL procedure and shares its parameters with collaborators. The client uses predictions from another participant’s randomly selected organ-specific model during the local training.